In a cartoon , a malfunctioning cannon fires a hungry coyote towards the bottom of a cliff with an initial rate of 100 feet per second. If the cliff is 1250 feet tall, how long will it take the coyote to reach the desert floor? ( To account for gravity , use the formula d - rt + 16^2, where d = dustance, r = initiall rate, and t = time).
Answer by stanbon(48558)

In a cartoon , a malfunctioning cannon fires a hungry coyote towards the bottom of a cliff with an initial rate of 100 feet per second. If the cliff is 1250 feet tall, how long will it take the coyote to reach the desert floor?
-------------------------
-16t^2-100t = -1250
16t^2 + 100t - 1250 = 0
(4t-25)(4t+50) = 0
Positive solution:
t = 25/4 = 6.25 seconds
===========================
Cheers,
Stan H.